# Statistical Taylor Expansion111

### v8

@@ 1.3.1 Conventional Floating-Point Arithmetic
$64919121\times 205117922-159018721\times 83739041=\\ 64919121.000000000\times 205117922.000000000-159018721.000000000\times 83739041.000000000=\\ 13316075197586562.-13316075197586560.=2.=2.0000000000000000;$ (1.2)
@@
    The normalization of the input values pads zero tail bit values artificially, to make each value as precise as possible for the 64-bit IEEE floating-point representation, e.g., from $64919121$ to $64919121.000000000$.
@@ 1.4 An Overview of This Paper
This paper presents the theory for statistical Taylor expansion, the implementation as variance arithmetic, and the tests of variance arithmetic. Section 1 compares statistical Taylor expansion and variance arithmetic with each known uncertainty-bearing arithmetic. Section 2 presents the theoretical foundation for statistical Taylor expansion. Section 3 presents variance arithmetic. Section 4 discusses the standards and methods to validate variance arithmetic. Section 5 tests variance arithmetic for the most common math library functions. Section 6 applies variance arithmetic to adjugate matrix and matrix inversion. Section 7 applies variance arithmetic to process time-series data. Section 8 focuses on the effect of numerical library errors and shows that such library errors can be significant. Section 9 applies variance arithmetic to regression. Section 10 provides a summary and discussion for variance arithmetic.
@@ 2.1 The Uncorrelated Uncertainty Assumption
The uncorrelated uncertainty assumption [1] states that the uncertainty of any input is statistically uncorrelated with any other input uncertainty. It is satisfied when uncorrelated noises are the major contribution of the uncertainty of the signals. It is consistent with the common methods in processing experimental data [2][3], such as the common knowledge or belief that precision improves with the count $n$ as $1/\sqrt{n}$ during averaging.
@@ 2.6 Convergence
2.7 Bounding Factor and Sample Count
[…]
figure The measured bounding factor $\kappa_{r}$ (as shown by the y-axis on the left) and the corresponding measured bounding leakage $\epsilon(\kappa_{r})$ (as shown by the y-axis on the right) for the different sample count $N$ (as shown by the x-axis) when the underlying distribution is either uniform or normal (as shown by the legend, with different measuring bounding factor for the normal distribution).
[…]
Generally, $\kappa$ in Formula (2.4) increases with the sample count $N$, which has to be sufficiently large to get accurate result using statistical Taylor expansion.
2.8 Approximation for Insufficient Sample Count
[…]
figure The result uncertainty or precision (as shown by the y-axis on the left) and the bounding leakage (as shown by the y-axis on the right) for the different bounding factor $\kappa$ (as shown by the x-axis) for $(1\pm 0.05)^{-1}$, $(1\pm 0.1)^{-1}$, $(1\pm 0.1)^{2}$, $\log(1\pm 0.1)$, $e^{1\pm 0.1}$ and $\sin(0\pm 0.1)$. For better comparison, the precision of $(1\pm 0.05)^{-1}$ is normalized by $1/2$ while the precision of $(1\pm 0.1)^{2}$ is normalized by $2$.
[…]
figure The variance error (as shown by the y-axis) after normalized by $\zeta(2,\kappa)$ for the different bounding factor $\kappa$ (as shown by the x-axis) for $(1\pm 0.05)^{-1}$, $(1\pm 0.1)^{-1}$, $(1\pm 0.1)^{2}$, $\log(1\pm 0.1)$, $e^{1\pm 0.1}$ and $\sin(0\pm 0.1)$ (as shown by the legend).
When $N$ or $\kappa$ is not sufficiently large, the result variance of statistical Taylor may be too small due to too much bounding leakage. Figure 2.8 shows that the result precision of $(1\pm 0.05)^{-1}$, $(1\pm 0.1)^{-1}$, $(1\pm 0.1)^{2}$, and $\log(1\pm 0.1)$, and the result uncertainty of $e^{1\pm 0.1}$ and $\sin(0\pm 0.1)$, are all less than the corresponding stable values when $\kappa=5$ for an input with Gaussian uncertainty. To compensate for this effect, the result can be normalized by $\zeta(2,\kappa)$. Figure 2.8 shows that the error after such normalization is much less than those of Figure 2.8 (e.g., from about 70% to less than 0.5% for $x^{2}$ when $\kappa=1$), although the results of such adjustment vary for different functions (e.g., smaller for $\sin(x)$ than $\log(x)$), different parameters (e.g., much less for $x^{2}$ than $x^{-1}$), and different input uncertainty (e.g., less for $(1\pm 0.05)^{-1}$ than $(1\pm 0.1)^{-1}$), so the result of such normalization needs to be analyzed case by case.
On the other hand, either $\log(x\pm\delta x)$ or $(x\pm\delta x)^{c}$ will not converge if the input $P(x)$ is larger than $1/\kappa$. To obtain the result, an artificial $\kappa<1/P(x)$ can be used then normalized to obtain an approximate result.
@@ 3 Variance Arithmetic
Variance arithmetic implements statistical Taylor expansion, by adding practical assumptions.
3.1 Choices of the Bounding Factor $\kappa$
@@
3.2 Low-Order Approximation
When $\kappa=5,n<5$, $\eta(n)\simeq n!!$!. Formula (2.27), (2.33), (2.30), and (2.36) can be simplified as Formula (3.1), (3.2), (3.2), (3.2), respectively.
$+\frac{7}{6}c^{2}(c-1)(c-2)^{2}(c-\frac{16}{7})P(x)^{6}+o(P(x)^{8});$ (3.4)
Formula (3.5), (3.6), and (3.7) are special cases for Formula (3.2).
$\delta^{2}x^{2}\simeq$ $4x^{2}(\delta x)^{2}+2(\delta x)^{4};$ (3.5)
3.3 Expansion Calculation
@@
[…]
figure The measured upper bound $\delta x$ for $1/(1\pm\delta x)$ for different $\delta y$ of Taylor coefficient uncertainty as shown by the x-axis, using the y-axis on the left. Also shown are the corresponding uncertainty bias and uncertainty, using the y-axis on the right.
Due to digital limitation of conventional floating-point representation, $\zeta(2n)$ can only be calculated to $4422n<$ 442 before becoming infinitive, Thus, variance arithmetic has different convergence rule from its counterpart of Taylor expansion of precise variable. In variance arithmetic, an infinitive expansion (including Taylor expansion) can fail for the following reasons: not monotonic, not stable, not positive, not finite, and not reliable.
3.3.1 Convergence
[…]
According to Formula (2.6), the condition for Formula (2.33) for $\log(x\pm\delta x)$ to converge is $P(x)\leq 1/\kappa=1/5$, which is confirmed numerically by $\log(x\pm\delta x)$ as $P(x)\lesssim 0.20088$. Beyond the upper bound $P(x)$, the expansion is not monotonic. Variance arithmetic rejects the distributional zero of $\log(0)$ inside the range of $[x-\delta x,x+\delta x]$ statistically due to the divergence of Formula (2.33) mathematically, with $\zeta(2n)$ as the connection of these two worlds.
According to Formula (2.38), except when c is a nature number, Formula (2.36) for $(x\pm\delta x)^{c}$ converges near $P(x)\simeq 1/\kappa=1/5$, with the upper bound $P(x)$ increasing with $c$. This is confirmed approximately by Figure 3.3. Beyond the upper bound $P(x)$, the expansion is not monotonic. Qualitatively, $\delta^{2}1/x$ converges slower than $\delta^{2}\sqrt{x}$ according to Formula (3.7) and (3.6).
3.3.2 Stable
Conventional condition to terminate an infinitive series numerically is when the residual of an expansion is less than the unit in the last place of the result [13]. In practice, the last expansion term usually proxies for the expansion residual if either the residual is known to be less than the last expansion term, or if the residual is difficult to calculate [13]. This method no longer works in variance arithmetic because the expansion is limited to $4422n<$ 442, and because the expansion result is no longer precise. The probability of an imprecie value $x\pm\delta x$ not equal 00 is indicated by its z-statistic $|x|/\delta x$ [5]. If this probability is the bounding leakage $\epsilon$, Formula (3.8) gives the corresponding upper bound $\tau$ for this z-statistic. Because variance arithmetic has already guaranteed that the expansion is absolutely and monotonically decreasing, if the absolute value of the last expansion term is more than $\tau$-fold of the result uncertainty, the result is invalidated as not stable.
$\xi(\frac{\tau}{\sqrt{2}})=\epsilon=1-\xi(\frac{\kappa}{\sqrt{2}}):\;\;\;\tau\simeq\epsilon\frac{\sqrt{2\pi}}{2}=7.18\times 10^{-7};$ (3.8)
3.3.3 Valid
At any order, if the expansion variance becomes negative, the result is invalidated as not positive.
The variance expansion can contain negative values, such as Formula (2.30) for $sin(x\pm\delta x)$. Figure 3.3 shows that the upper bound $\delta x$ for $sin(x\pm\delta x)$ is periodic between $0.318\pi$ and $0.416\pi$, which is nicely less than the periodicity $2\pi$ of $\sin(x)$. Beyond the upper bound $\delta x$, the expansion is invalidated as not positive in Figure 3.3.
3.3.4 Finite
If the result of Taylor expansion is not finite, the expansion is invalidated as not finite. This usually happens when the expansion diverges so that it is usually not the reason to bound $\delta x$. For example $(1\pm 1)^{-2}$ has infinitive variance.
@@ 3.3.5 Reliable
While $\delta x$ specifies the uncertainty of $x$, the calculated $\delta x$ itself can be imprecise. To guarantee $\delta x$ to be precise enough, during expansion, if the value of the variance is less than the $1/\kappa$-fold of its uncertainty, the result is invalidated as not reliable.
Fortunately, being not reliable is rarely the reason to bound $\delta x$.
3.3.6 Imprecise Taylor Coefficients
[…]
3.3.7 Continuous
@@
For $x^{c}$, when $c$ is a nature number $n$, Formula (2.36) has $n$ terms, so that $\delta x$ has no upper bound $\delta x$ for $(x\pm\delta x)^{n}$. The result mean, variance and histogram of $(x\pm\delta x)^{c}$ are continuous around $c=n$. Table 3.3.7 shows that the result of $(1\pm\delta x)^{n\pm d}$ is very close to that of $(1\pm\delta x)^{n}$, even the former had upper bound for $\delta x$, while the latter has not.
@@
A statistical bounding range of variance arithmetic cannot include more than one distributional pole because the corresponding statistical Taylor expansion becomes not positive. One such example is $\sin(x)$ in Figure 3.3.
@@
3.4 Numerical Representation
When the uncertainty $\delta x$ is not specified initially:
[…]
@@ 3.5 Tracking Rounding Error
figure The residual error of $\sum_{j=0}^{121}x^{j}-\frac{1}{1-x}$ vs $x$. The y-axis to the left shows the value and uncertainty of residual errors. The uncertainty calculated by $\sum_{j=0}^{121}x^{j}$ overlaps with that by $\frac{1}{1-x}$. The y-axis to the right shows the expansion order needed to reach stable value for each $x$.
@@
Figure 3.5 shows the residual error of $\sum_{j=0}^{121}x^{j}-\frac{1}{1-x}$, in which the polynomial $\sum_{j=0}^{121}x^{j}$ is calculated using Formula (3.9), $\frac{1}{1-x}$ is calculated using Formula (2.34), and $x$ is initiated as a floating-point value. Because $\eta(2n)$ is limited to $2n\leq 242$, Formula (3.9) for polynomial is limited to $N\leq 121$, so that $\sum_{j=0}^{121}x^{j}$ has lower expansion order than that of $\frac{1}{1-x}$. Figure 3.5 shows:
  • When $x\in[-0.73,0.75]$, the required expansion order is no more than $121$, so that the residual error is the calculation rounding error between $\sum_{j=0}^{121}x^{j}$ and $\frac{1}{1-x}$. Detailed analysis shows that the max residual error is $4$-fold of the $LSV$ of $\frac{1}{1-x}$. The calculated uncertainty bounds the residual error nicely for all $x\in[-0.73,0.75]$.
  • When $x\not\in[-0.74,+0.75]$, such as when $x=-0.75,0.76$, the required expansion order is more than $121$, so that the residual error is caused by insufficient expansion order of $\sum_{j=0}^{121}x^{j}$, and the residual error increases absolutely when $|x|\rightarrow 1$, e.g., reaching $50$ when $x=0.98$.
@@ 5 Math Library Functions
figure The values and uncertainties of $\log(e^{x})-x$ vs $x$, as VarDbl Error and VarDbl Uncertainty in the legend. The result of the same calculation using conventional floating-point library functions is shown as Lib Error in the legend.
@@ 5.5 Numerical Errors for Library Functions
The combined numerical error of the library function $e^{x}$ and $\log(x)$ is calculated as $\log(e^{x})-x$. Figure 5 shows that using either variance arithmetic or the conventional floating-point library functions results in the same value errors. The uncertainties of variance arithmetic bound the value errors effectively, resulting in an error deviation about $0.409$ when $|x|\leq 1$. It looks suspicious when $|x|>1$ the value error is always 00 in both calculations.

### v9

@@ 1.3.1 Conventional Floating-Point Arithmetic
$64919224\times 205117922-159018721\times 83739041=\\ 64919224.000000000\times 205117922.000000000-159018721.000000000\times 83739041.000000000=\\ 13316075197586562.-13316075197586560.=2.=2.0000000000000000;$ (1.2)
@@
    The normalization of the input values pads zero tail bit values artificially, to make each value as precise as possible for the 64-bit IEEE floating-point representation, e.g., from $64919224$ to $64919224.000000000$.
@@ 1.4 An Overview of This Paper
This paper presents the theory for statistical Taylor expansion, the implementation as variance arithmetic, and the tests of variance arithmetic. Section 1 compares statistical Taylor expansion and variance arithmetic with each known uncertainty-bearing arithmetic. Section 2 presents the theoretical foundation for statistical Taylor expansion. Section 3 presents the variance arithmetic as an implementation of statistical Taylor expansion. Section 4 discusses the standards and methods to validate variance arithmetic. Section 5 tests variance arithmetic for the most common math library functions. Section 6 applies variance arithmetic to adjugate matrix and matrix inversion. Section 7 applies variance arithmetic to process time-series data. Section 8 focuses on the effect of numerical library errors and shows that such library errors can be significant. Section 9 applies variance arithmetic to regression. Section 10 provides a summary and discussion for variance arithmetic.
@@ 2.1 The Uncorrelated Uncertainty Assumption
The uncorrelated uncertainty assumption [1] states that the uncertainty of any input is statistically uncorrelated with any other input uncertainty. It is satisfied when uncorrelated noises are the major contribution of the uncertainty of the signals. It is consistent with the common methods in processing experimental data [2][3].
@@ 2.3 Statistical Taylor Expansion
2.4 Bounding Factor and Sample Count
[…]
figure The measured bounding factor $\kappa$ (as shown by the y-axis on the left) and the corresponding measured bounding leakage $\epsilon(\kappa)$ (as shown by the y-axis on the right) for the different sample count $N$ (as shown by the x-axis) when the underlying distribution is either uniform or normal (as shown by the legend, with different measuring bounding factor for the normal distribution).
[…]
figure The result variance ratio $\delta^{2}f/\widehat{\delta^{2}f}$ (as shown by the y-axis on the left) and the bounding leakage (as shown by the y-axis on the right) for the different bounding factor $\kappa$ (as shown by the x-axis) for the selected $f(x)$ (as shown by the legend), in which $\widehat{\delta^{2}f}$ is the corresponding variance for $\kappa=5$.
[…]
figure The ratio of adjusted $\widehat{\delta^{2}f}$ to the original $\widehat{\delta^{2}f}$ (as shown by the y-axis) for the different bounding factor $\kappa$ (as shown by the x-axis) for the selected $f(x)$ (as shown by the legend), in which $\widehat{\delta^{2}f}$ is approximated by normalizing the input variance with $\zeta(2,\kappa)$ so that $\delta^{2}x=(\delta x)^{2}$.
@@ 2.4 Bounding Factor and Sample Count
[…]
Generally, $\kappa$ in Formula (2.4) increases with the sample count $N$, which has to be sufficiently large to get accurate result using statistical Taylor expansion. Let $\widehat{\delta^{2}f}$ be the ideal variance when the result variance becomes stable with increasing $\kappa$. Figure 2.4 shows that $\delta^{2}f$ increases with $\kappa$ until reaches $\widehat{\delta^{2}f}$ for the selected functions. The result bounding factor is defined by Formula (2.13) using the result variance ratio, e.g., Formula (2.14) shows that the result $\kappa$ for $x\pm y$ is a variance weighted average of the input $\kappa(N_{x})$ and $\kappa(N_{y})$ through $\zeta(2,\kappa)$. Figure 2.4 also shows that $\kappa$ remains roughly a constant during calculation, with all the result variance ratios very close to that of $x$ for all the input $\kappa$.
$\zeta(2,\kappa)\equiv$ $\;\frac{\delta^{2}f}{\widehat{\delta^{2}f}};$ (2.13)
$x\pm y:$ $\;\;\;\zeta(2,\kappa)=\frac{\zeta(2,\kappa(N_{x}))(\delta x)^{2}+\zeta(2,\kappa(N_{y}))(\delta y)^{2}}{(\delta x)^{2}+(\delta y)^{2}};$ (2.14)
The cutoff $\kappa$ for $\widehat{\delta^{2}f}$ is somewhat subjective. In this paper, it is when $\kappa$ reaches $5$ or the input sample count $N$ reaches $1000$ for Gaussian distribution.
As stated later, when $1/5<P(x)$, $\log(x\pm\delta x)$ or $(x\pm\delta x)^{c}$ diverges for $\widehat{\delta^{2}f}$. To approximate $\widehat{\delta^{2}f}$, an artificial $\kappa_{a}$ satisfying $P(x)^{2}\leq\zeta(2,\kappa_{a})/\kappa_{a}^{2}$ can be used to normalize the input variance $(\delta x)^{2}$ as $(\delta x)^{2}/\zeta(2,\kappa_{a})$, to achieve $\delta^{2}x=(\delta x)^{2}$ perfectly for all $\kappa$. Figure 2.4 shows that the result variances are close to those of $5<\kappa$, although as a first order approximation, the results depend on the actual $f(x)$. Figure 2.4 shows a universal worst $\kappa\simeq 1.75$, below which the tail effect of Gaussian distributions disappears.
@@ 2.6 One Dimensional Examples
2.7 Low-Order Approximation
When $n<5,5\leq\kappa$, $\eta(n)\simeq n!!$!. Formula (2.31), (2.37), (2.34), and (2.40) can be simplified as Formula (2.41), (2.42), (2.7), (2.7), respectively.
$+\frac{7}{6}c^{2}(c-1)(c-2)^{2}(c-\frac{16}{7})P(x)^{6}+o(P(x)^{8});$ (2.44)
Formula (2.45), (2.46), and (2.47) are special cases for Formula (2.7).
$\delta^{2}x^{2}\simeq$  $\;4x^{2}(\delta x)^{2}+2(\delta x)^{4};$ (2.45)
@@ 2.9 Dependency Tracing
  • If no parameters $x_{1},x_{2},...$ of $g(x_{1},x_{2},...)$ appears directly in $f(g,y_{1},y_{2},..)$, then $\widehat{\delta^{2}g}$ and the corresponding $\zeta(2,\kappa)$ can be used as the direct input to $f$ to calculate $\widehat{\delta^{2}f}$ and $\delta^{2}f$. Otherwise, the additional dependency between $f$ and $g$ through $x_{1},x_{2},...$ can not be ignored in general.
@@ 3 Variance Arithmetic
This paper limits to calculating the variances for sufficient sample count $\delta^{2}f$, or ideal variances $\widehat{\delta^{2}f}$ when $\kappa=5$, and the input distribution is limited to Gaussian.
3.1 Numerical Representation
When the uncertainty $\delta x$ is not specified initially:
@@ 3.1 Numerical Representation
3.2 Calculation Rules
Variance arithmetic implements statistical Taylor by adding practical assumptions. Due to digital limitation of conventional floating-point representation, $\zeta(2n)$ can only be calculated to $4462n<$ 446 before becoming infinitive, so practical numerical rules are introduced for:
  • finite: The result should be finite.
[…]
  • stable: For the convergence to be fast enough, the value of the last expansion term should be less than the $\epsilon(5)$-fold of the result uncertainty and the result value.
  • positive: At any order, the expansion variance has to be positive.
  • reliable: At any order, the uncertainty of the variance should be less than $1/\kappa$-fold of the uncertainty of the variance.
The validity of the above numerical rules of variance arithmetic will be demonstrated in the following.
3.2.1 Monotonic
@@
According to Formula (2.8), the condition for Formula (2.37) for $\log(x\pm\delta x)$ to converge is $P(x)\leq 1/\kappa=1/5$, which is confirmed numerically by $\log(x\pm\delta x)$ as $P(x)\lesssim 0.20087$. Beyond the upper bound $P(x)$, the expansion is not monotonic. Variance arithmetic rejects the distributional zero of $\log(0)$ inside the range of $[x-\delta x,x+\delta x]$ statistically due to the divergence of Formula (2.37) mathematically, with $\zeta(2n)$ as the connection of these two worlds.
According to Formula (2.49), except when c is a nature number, Formula (2.40) for $(x\pm\delta x)^{c}$ converges near $P(x)\simeq 1/\kappa=1/5$, with the upper bound $P(x)$ increasing with $c$. This is confirmed approximately by Figure 3.2.1. Beyond the upper bound $P(x)$, the expansion is not monotonic. Qualitatively, $\delta^{2}1/x$ converges slower than $\delta^{2}\sqrt{x}$ according to Formula (2.47) and (2.46).
3.2.2 Stable
The condition for not stable always happens together with the not monotonic condition, but seldom happens by itself.
3.2.3 Positive
@@
The variance expansion can contain negative values, such as Formula (2.34) for $sin(x\pm\delta x)$. Figure 3.2.3 shows that the upper bound $\delta x$ for $sin(x\pm\delta x)$ is periodic between $0.318\pi$ and $0.416\pi$, which is nicely less than the periodicity $2\pi$ of $\sin(x)$. Beyond the upper bound $\delta x$, the expansion is invalidated as not positive in Figure 3.2.3.
@@ 3.2.4 Reliable
The condition for not reliable seldom happens.
3.3 Continuity
@@
For $x^{c}$, when $c$ is a nature number $n$, Formula (2.40) has $n$ terms, so that $\delta x$ has no upper bound $\delta x$ for $(x\pm\delta x)^{n}$. The result mean, variance and histogram of $(x\pm\delta x)^{c}$ are continuous around $c=n$. Table 3.3 shows that the result of $(1\pm\delta x)^{n\pm d}$ is very close to that of $(1\pm\delta x)^{n}$, even the former had upper bound for $\delta x$, while the latter has not.
@@
A statistical bounding range of variance arithmetic cannot include more than one distributional pole because the corresponding statistical Taylor expansion becomes not positive. One such example is $\sin(x)$ in Figure 3.2.3.
@@ 3.4 Tracking Rounding Error
figure The residual error of $\sum_{j=0}^{224}x^{j}-\frac{1}{1-x}$ vs $x$. The y-axis to the left shows the value and uncertainty of residual errors. The uncertainty calculated by $\sum_{j=0}^{224}x^{j}$ overlaps with that by $\frac{1}{1-x}$. The y-axis to the right shows the expansion order needed to reach stable value for each $x$.
@@
Figure 3.4 shows the residual error of $\sum_{j=0}^{224}x^{j}-\frac{1}{1-x}$, in which the polynomial $\sum_{j=0}^{224}x^{j}$ is calculated using Formula (3.1), $\frac{1}{1-x}$ is calculated using Formula (2.38), and $x$ is initiated as a floating-point value. Because $\eta(2n)$ is limited to $2n\leq 242$, Formula (3.1) for polynomial is limited to $N\leq 224$, so that $\sum_{j=0}^{224}x^{j}$ has lower expansion order than that of $\frac{1}{1-x}$. Figure 3.4 shows:
  • When $x\in[-0.73,0.75]$, the required expansion order is no more than $224$, so that the residual error is the calculation rounding error between $\sum_{j=0}^{224}x^{j}$ and $\frac{1}{1-x}$. Detailed analysis shows that the max residual error is $4$-fold of the $LSV$ of $\frac{1}{1-x}$. The calculated uncertainty bounds the residual error nicely for all $x\in[-0.73,0.75]$.
  • When $x\not\in[-0.74,+0.75]$, such as when $x=-0.75,0.76$, the required expansion order is more than $224$, so that the residual error is caused by insufficient expansion order of $\sum_{j=0}^{224}x^{j}$, and the residual error increases absolutely when $|x|\rightarrow 1$, e.g., reaching $50$ when $x=0.98$.
@@ 5 Math Library Functions
figure The values and uncertainties of $\log(e^{x})-x$ and $e^{\log(x)}-x$ vs $x$, with $0.1$ as $x$ intervals.
@@ 5.5 Numerical Errors for Library Functions
The combined numerical error of the library function $e^{x}$ and $\log(x)$ is calculated as either $\log(e^{x})-x$ or $e^{\log(x)}-x$. Using either variance arithmetic or the conventional floating-point library functions results in the same value errors. Figure 5 shows the corresponding result uncertainties, which has a minimal value when the input is a 2’s fractional such as $1$ or $1/2$. The normalized error is $0.409$ when $|x|\leq 1$ for $\log(e^{x})-x$, and almost zero otherwise.